The ship returned to Hole 858G in DP mode during the trip to pickup the modified CORK running tool. The environmental conditions were good (seas 2 ft, swells 5 ft, heave 0.6 m). The hole was reentered in 15 min; however, the running tool rotated freely indicating that it had not jayed-in. After putting 15 K lb on the running tool, an overpull to 10 K lb indicated successful jay-in. The data logger overshot was run on wireline, and an unsuccessful attempt was made to retrieve the data logger by jarring and pumping with 15 K lb weight down to 10 K lb tension on the running tool. The running tool appeared to release easily, and the TV showed the upper portion of the CORK and possibly the upper mandrel and data logger (without instrument cable) engaged in the running tool body. The overshot was retrieved, and the running tool was retrieved; however, the fish was lost. The jay-slots in the running tool had engaged the CORK; therefore, they must have un-jayed and dropped the fish while pulling back to the ship.

The CORK pulling tool was run in an attempt to retrieve the bottom latch and seal portion of the CORK. The reentry required 2.25 hr and was more difficult because the flat skirt on the tool was difficult to position in the small throat on the flat ROV platform. A tapered throat on the ROV platform would aid in reentry. The CORK pulling tool was spudded and circulated down several meters into the throat with 15 Kips and 250 gpm at 150 psi; however, the tool would not jay into the lower CORK body. The evidence indicated that the running tool might be hanging up on the ROV platform. The ROV platform pulling tool was modified to change it from wireline to free fall deployment. It was dropped and engaged the ROV platform grating with both arms.

A considerable quantity of material was circulated out of the reentry cone throat. The pipe had been worked up and down and circulated while installing the platform tool, and the pipe had gone down another 1.5 m over the CORK. When an attempt was made to pull the platform pulling tool, the CORK tool was found to have been jayed-in while washing and working the tool. The CORK was pulled with increasing tension and circulated for 45 min., finally coming free with 100,000 lb overpull.

The lower CORK body and ROV platform were pulled to the ship in 4.25 hr, and the tools were laid down in 5 hr, including sampling for biological samples and minerals. The ROV platform was only held by one blade on the platform pulling tool, and both blades on the other arm had been overloaded and bent down. The inside of the lower CORK body was full of 3" thick mineralized deposits. The viton seals were brittle and had failed, allowing hydrothermal fluids to circulated past the seals and through the upper CORK body until the mineral deposits choked off flow. The nuts were gone (by cathodic corrosion?) from 4 of 8 release rods. All surfaces were coated with fine black film about 1 mm thick. The 400-m-long kevlar thermistor string and 2-m-long stainless steel weight bar remained in the hole. No trace of the kevlar was found.

A 9-7/8" RBI C-7 RCB bit was run to clean out the hole if required, and the bit was set at 2446 mbrf (20 mbsf) for fluid sampling. The WSTP tool was run to 2443 m, and a good hydrothermal fluid sample was obtained. However, the 150°C limit of the temperature tool was approached, the gradient record was lost, and an O-ring leak shorted out the electronics. The "Ultra High Temperature-Multi Sensor Memory-Geophysical Research Corporation" (UHT-MSM-GRC or Becker) temperature tool was run for the first time ever. The tool is 8.5 ft long x 2-1/8 in. OD, has a temperature rating of 350°C, and is run on the coring wireline. The Becker tool set down on an obstruction at 2631 mbrf (205 mbsf = 64.7 m above the 11-3/4" casing shoe) and read a constant 273°C from 205 to 20 m, which indicates flow up the hole. The lithium batteries were 90°C when the tool was retrieved; therefore, the batteries were allowed to cool before extracting the samples. The lithium batteries have a temperature rating of 150°C and could explode at 180°C. A safety procedure was adopted to cool the batteries on future runs.

The Los Alamos fluid sampler tool was run in to 200 m (2526 mbrf); however, the viton valve seats melted and the hydrothermal fluid sample had mixed with seawater on the wireline trip out. The bit tagged fill at 205 m (same as the Becker tool). To avoid circulating cold seawater into the hole, an unsuccessful attempt was made to rotate down without circulation. The hole was cleaned out to 387 m in soft material, but the pipe stuck at 326 mbsf when pulling out of the hole. The pipe was worked for 4 hr at up to 200 K lb overpull, the rotary was stalled at 800 amps, and the hole was packing off (500 psi at 50 gpm). To avoid contaminating the hole for water samples, no mud had been pumped while cleaning out the hole as requested. The kevlar thermistor string was initially suspected as causing the stuck pipe incident by wadding up around the BHA; however, an inspection with the bit in the reentry cone showed the BHA was clean. A 20 and 30 bbl tandem sepiolite mud sweep was circulated and immediately freed the pipe.

Wash barrel 169-858G-18W was retrieved with loose sulfides, anhydrites, and plastic electrical wrap and insulation from the thermistor string. No kevlar was ever recovered, leading to speculation that it had fallen apart. The problem was evidently a combination of a boulder or ledge sticking at the shoe and inadequate hole cleaning with seawater. The pipe actually stuck a second time at 278 mbsf because circulation stopped before the mud cleared the seafloor. The reentry cone throat in the CORK setting and seal area was washed and rotated in with the bit, and a short trip was made to 385 mbsf without tagging fill. The 9-7/8" RBI CC-7 bit was pulled after 31 hr in 273°C hydrothermal fluid. The body was blackened, the cones rotated easily, and a few teeth were chipped; however, the bit still appeared usable.

The CORK BHA was run as follows: stinger of 2 joints of 7-1/4" DC, crossover, CORK assembly (11-3/4" stretch version), CORK setting tool, seal bore pony DC, 5 joints 8-1/4" DC, tapered DC, 6 joints of 5-1/2" DP, crossover. The hole was reentered at 2130 hr on 4 September. A 400-m-long kevlar-encased thermistor string (8 sensors) was run with a 2-m stainless steel sinker bar on bottom. The data logger was latched in and had to be jarred off. The CORK was landed and latched in at 1000 psi. A 10K lb pull verified latch in. The CORK was released at 0745 hr on 5 September. A new solid deck ROV platform was dropped, and a TV inspection showed the platform and CORK were in good shape. The beacon was turned off for a later return to the Dead Dog sites. The weather forecast was good for the next few days so a decision was made to go to Hole 857D for another CORK replacement job.

To 169 Hole 857D

To 169 Table of Contents

Publications Homepage

ODP Homepage